# DocForge-Bench: A Comprehensive Benchmark for Document Forgery Detection and Analysis

### v1

@@ Abstract
We present DocForge-Bench, the first unified zero-shot benchmark for document forgery detection, evaluating 14 methods across eight datasets spanning text tampering, receipt forgery, and identity document manipulation. Unlike fine-tuning-oriented evaluations such as ForensicHub (Du and others, 2025), DocForge-Bench applies all methods with their published pretrained weights and no domain adaptation—a deliberate design choice that reflects the realistic deployment scenario where practitioners lack labeled document training data. Our central finding is a pervasive calibration failure invisible under single-threshold protocols: methods achieve moderate Pixel-AUC (0.76) yet near-zero Pixel-F1. This AUC–F1 gap is not a discrimination failure but a score-distribution shift: tampered regions occupy only 0.27–4.17% of pixels in document images—an order of magnitude less than in natural image benchmarks—making the standard threshold catastrophically miscalibrated. Oracle-F1 is 2–10 higher than fixed-threshold Pixel-F1, confirming that calibration, not representation, is the bottleneck. A controlled calibration experiment validates this: adapting a single threshold on domain images recovers 39–55% of the Oracle-F1 gap, demonstrating that threshold adaptation—not retraining—is the key missing step for practical deployment. Overall, no evaluated method works reliably out-of-the-box on diverse document types, underscoring that document forgery detection remains an unsolved problem. We further note that all eight datasets predate the era of generative AI editing; benchmarks covering diffusion- and LLM-based document forgeries represent a critical open gap on the modern attack surface.
@@ 1 Introduction
Extreme region imbalance: Document forgeries typically modify a few characters or fields, leaving 95–99.7% of pixels authentic. Natural image benchmarks assume 10–30% tampered area; document datasets range from 0.27% (ReceiptForgery) to 4.17% (FSTS-1.5k). This order-of-magnitude difference invalidates the standard decision threshold used by all published methods.
@@
Mechanistic explanation of the gap: We show quantitatively that the AUC–F1 gap is driven by tampered-pixel base rates (0.27–4.17%) that are 10–50 lower than in natural image benchmarks, making catastrophically miscalibrated—and that this is correctable via threshold adaptation on a small domain sample, without retraining.
@@ 2.2 Document-Specific Forgery Detection
VLM-based approaches. Recent work fine-tunes large vision-language models for document forgery, including TextShield-R1 (Qu and others, 2026) and LogicLens (Zeng et al., 2025). These represent a promising direction but are outside the scope of DocForge-Bench, which focuses on pixel-level localization methods with frozen pretrained weights. Liang and others (2025) find that even frontier LLMs struggle with pixel-level localization—a finding complementary to ours.
@@ 3.1 Text-Region Tampering
We evaluate this threat across six complementary datasets spanning different document domains and imaging conditions: DocTamper (Qu et al., 2023a) provides large-scale synthetic diversity (170K images) across document layouts. T-SROIE (Zhang and others, 2022) captures receipt-context tampering with realistic OCR backgrounds. RealTextManipulation (Liao and others, 2022) provides authentic forgeries collected in the wild, without synthesis artifacts. Tampered-IC13 (Wang et al., 2022b) covers in-the-wild scene-text images (storefronts, signage) rather than document scans. MixTamper (Li et al., 2025) introduces multi-label tampering categories (copy-move, splicing, text generation, smearing, erasing). FSTS-1.5k (Yu et al., 2025) provides a real-world evaluation derived from 16,750 human-annotated forgery instances, capturing the authentic distribution of text forgery parameters.
@@ 3.4 Scope and Coverage
Table 1 maps each dataset to the threat scenarios above and summarises evaluation coverage. Together, these eight datasets span four dimensions of variation critical for robust evaluation: (1) Realism: synthetic (DocTamper, MixTamper) vs. real-world (RealTextManipulation, ReceiptForgery, FSTS-1.5k); (2) Annotation: pixel-level masks vs. bounding boxes (ReceiptForgery, Tampered-IC13); (3) Document type: formal documents, receipts, scene text, and identity cards; (4) Scale: from 218 images (ReceiptForgery) to 170K (DocTamper).
We do not evaluate AI-generated document content (GANs, diffusion models), LLM-generated text detection, or physical print-scan-reprint attacks in this benchmark. These represent important adjacent challenges but require distinct methods and evaluation protocols beyond the pixel-level localization framework used here.
@@ 4.1 Document-Specific Tampering Datasets
(Liao and others, 2022) provides 9,000 real-world text manipulation images collected from the internet, with pixel-level semantic segmentation masks. Unlike synthetically generated benchmarks, this dataset captures authentic forgeries with diverse backgrounds, text styles, and manipulation strategies. The official test split (3,197 images) contains 1,203 tampered images and 1,994 authentic images (“good_*” prefix, empty masks); we evaluate on the tampered-only subset (1,000 sampled after exclusion of authentic images).
@@
(Mesquita et al., 2023) (L3i / ICDAR 2023 Competition) contains 988 receipt photographs (577 train / 193 val / 218 test), of which 163 across all splits are forged; the test split contains 35 forged images out of 218 (16% positive rate). Forgeries were produced with common image editors (GIMP, Paint) by replacing printed price or quantity fields. Annotations are VGG-format rectangular bounding boxes embedded in a CSV manifest; pixel masks are rasterized from bounding boxes for localization evaluation. Of the 218 test images, 183 are authentic (empty masks) and are excluded from pixel-level metric computation via the NaN convention; metrics are aggregated over the 35 forged images only.
@@
(Li et al., 2025) is a multi-label document tampering dataset containing approximately 30,200 images with five tampering categories: copy-move, splicing, text-generating, smearing, and erasing. Ground truth is provided as RGB color-coded masks where each channel indicates a distinct tampering type; binary evaluation maps any non-zero pixel to tampered. Images are uniformly cropped to pixels and sourced from the StaVer and SCUT-EnsExam corpora. We use the 6,817-image test split and sample 1,000 for tractable evaluation.
@@
| Dataset | Domain | Year | Size | Forgery Types | GT Masks | Used In |
| DocTamper | Document | 2023 | 170K | Text replace/insert/delete | ✓ | Both |
| T-SROIE | Receipt | 2022 | 360 | Text tampering | ✓ | Both |
| RealTextManip. | Real-world | 2022 | 9K | Text manipulation | ✓ | Both |
| Tampered-IC13 | Scene text | 2015 | 233 | Text replacement | BBox∗ | Both |
| Receipt Forgery | Receipt | 2023 | 988 | Price/qty forgery | BBox∗ | Both |
| MixTamper | Document | 2024 | 30K | Copy-move, splice, gen., smear, erase | ✓(RGB) | Both |
| FSTS-1.5k | Real-world | 2025 | 1,488 | Text manipulation | ✓ | Both |
| FantasyID | Identity doc. | 2025 | 6,571 | Face swap, text replace | ✓ | Both |
@@ 5.2 Document-Specific Tampering Detection
All methods are evaluated with their official pretrained weights without fine-tuning. The general forensics methods (TruFor, ManTraNet, MVSS-Net, CAT-Net, PSCC-Net, IML-ViT, SAFIRE) were trained exclusively on natural photographic imagery—CASIA (Dong et al., 2013), FantasticReality (Kniaz et al., 2019), IMD2020 (Novozamsky et al., 2020), tampCOCO, RAISE, MS COCO, camera-model datasets (VISION, Dresden, KCMI), and proprietary synthetic databases. None saw any document images; all eight benchmark datasets are fully out-of-distribution for them. Among document-specific methods, DocTamper (model), DTD, FFDN, CAFTB-Net, and ADCD-Net were each trained on the DocTamperV1 training split (120,000 images)—their DocTamper test results are therefore in-domain, while all other datasets remain cross-domain. ASCFormer was trained on the RTM training split only (5,803 images); its RTM test results are in-domain. TIFDM was trained on a private, undisclosed corpus by the original authors (Dong et al., 2024) and is cross-domain on all eight datasets.
@@ 6.2 Experiments
We compare seven document-specific methods (DocTamper, DTD, FFDN, CAFTB-Net, TIFDM, ASCFormer, ADCD-Net) against all seven general methods (TruFor, ManTraNet, MVSS-Net, CAT-Net, PSCC-Net, IML-ViT, SAFIRE) on all eight document datasets. For datasets with bounding-box annotations only (ReceiptForgery, Tampered-IC13), predicted masks are evaluated against rasterized ground-truth boxes. MixTamper provides RGB color-coded multi-label masks; for binary evaluation any tampered channel maps the pixel to tampered. Authentic images (empty GT masks) are excluded from pixel-level metric aggregation via the NaN convention. This experiment directly measures the zero-shot advantage of domain-specific training and whether recent document forensics advances outperform both DocTamper and general forensics methods.
@@ 6.3 Implementation Details
All experiments are conducted using an NVIDIA GPU with at least 24 GB memory. Images are processed at each method’s native input resolution. For pixel-level evaluation, prediction masks are resized to match ground truth dimensions using bilinear interpolation. To maintain tractable evaluation on high-resolution datasets (T-SROIE receipt images can reach 49617016 px; FantasyID images average 32002000 px), both predictions and ground-truth masks are jointly downsampled to at most 2 megapixels using area interpolation before metric computation. This cap does not apply to images already within 2 MP. Evaluation scripts use scikit-learn for metric computation. For datasets exceeding the 1,000-image sample cap (DocTamper: 30K images; RealTextManipulation: 1,203 tampered images; MixTamper), images are selected using random.sample with seed 42, then sorted by filename for deterministic ordering. For ReceiptForgery, Pixel-IoU and Pixel-AUC are computed only over the 35 forged test images (16% of the 218-image test split); the remaining 183 authentic images produce NaN values under our convention and are excluded from the dataset mean.
@@
Methods vary in how they handle high-resolution inputs: TruFor processes images up to 40964096 px natively; MVSS-Net and PSCC-Net resize to their training resolution (512512 and 512512 respectively) before inference; IML-ViT pads/crops images to 10241024 for inference (only the top-left 10241024 region is evaluated on images larger than this; for FantasyID, an analysis of annotation bounding boxes shows that approximately 79% of forgery pixels fall outside this crop due to the distributed layout of identity card fields, so IML-ViT’s FantasyID score underestimates its true capability on full-card evaluation). The complete evaluation pipeline, including configs, scripts, and metric computation, is released as open-source.

### v2

@@ Abstract
We present DocForge-Bench, the first unified zero-shot benchmark for document forgery detection, evaluating 14 methods across eight datasets spanning text tampering, receipt forgery, and identity document manipulation. Unlike fine-tuning-oriented evaluations such as ForensicHub (Du and others, 2025), DocForge-Bench applies all methods with their published pretrained weights and no domain adaptation—a deliberate design choice that reflects the realistic deployment scenario where practitioners lack labeled document training data. Our central finding is a pervasive calibration failure invisible under single-threshold protocols: methods achieve moderate Pixel-AUC (0.76) yet near-zero Pixel-F1. This AUC–F1 gap is not a discrimination failure but a score-distribution shift: tampered regions occupy only 0.27–4.17% of pixels in document images—an order of magnitude less than in natural image benchmarks—making the standard threshold catastrophically miscalibrated. Oracle-F1 is 2–10 higher than fixed-threshold Pixel-F1, confirming that calibration, not representation, is the bottleneck. A controlled calibration experiment validates this: adapting a single threshold on domain images recovers 39–55% of the Oracle-F1 gap in representative evaluated high-AUC cases, demonstrating that threshold adaptation—not retraining—is the key missing step for practical deployment. Overall, no evaluated method works reliably out-of-the-box on diverse document types, underscoring that document forgery detection remains an unsolved problem. We further note that all eight datasets predate the era of generative AI editing; benchmarks covering diffusion- and LLM-based document forgeries represent a critical open gap on the modern attack surface.
@@ 1 Introduction
Extreme region imbalance: Document forgeries typically modify a few characters or fields, leaving 95–99.7% of pixels authentic. Natural image benchmarks assume 10–30% tampered area (Dong et al., 2013; Hsu and Chang, 2006; Wen et al., 2016); document datasets range from 0.27% (ReceiptForgery) to 4.17% (FSTS-1.5k). This order-of-magnitude difference invalidates the standard decision threshold used by all published methods.
@@
Mechanistic explanation of the gap: We show quantitatively that the AUC–F1 gap is driven by tampered-pixel base rates (0.27–4.17%) that are between 3 and 100 lower than in natural image benchmarks, making catastrophically miscalibrated—and that this is correctable via threshold adaptation on a small domain sample, without retraining.
@@ 2.2 Document-Specific Forgery Detection
VLM-based approaches. Recent work fine-tunes large vision-language models for document forgery, including TextShield-R1 (Qu and others, 2026) and LogicLens (Zeng et al., 2025). These represent a promising direction but are outside the scope of DocForge-Bench, which focuses on methods that produce dense pixel-level localization masks. VLM-based approaches typically produce natural language descriptions or coarse bounding regions rather than binary spatial masks; evaluating them under our pixel-F1/AUC protocol would require non-trivial post-processing that introduces design choices beyond our zero-shot evaluation scope (Liang and others, 2025). Pixel-level forensic localization with VLMs remains an open problem.
@@ 2.3 AI-Generated Content Detection
2.4 Evaluation Metrics and Protocols
Evaluation of forgery detection methods spans three distinct granularities, and the choice among them substantially affects reported numbers and inter-study comparability.
Pixel-level localization. The dominant paradigm in image manipulation detection treats forgery localization as binary segmentation: each pixel is labeled tampered or authentic, and performance is measured against a pixel-level ground-truth mask. The two primary metrics are Pixel-F1 and Pixel-AUC. Pixel-F1 at a fixed threshold is the harmonic mean of pixel-level precision and recall; it directly reflects deployment performance without calibration. Pixel-AUC is the area under the ROC curve swept over all thresholds and measures how well the model ranks tampered pixels above authentic ones, independent of calibration. A critical but underappreciated distinction is whether F1 is computed at a fixed threshold or at the per-image optimal threshold: the latter can inflate scores by 20–30 percentage points over fixed-threshold evaluation (SCU-ZJZ, 2024). This inconsistency pervades the document forgery literature—DTD (Qu et al., 2023b), FFDN (Chen et al., 2024), and ADCD-Net (Wong and others, 2025) all report F1 at an optimized threshold, while ForensicHub (Du and others, 2025) and IMDLBenCo (SCU-ZJZ, 2024) standardize on . Additional pixel-level metrics include IoU (Jaccard index; algebraically equivalent to ), Matthews Correlation Coefficient (MCC; balances all four confusion-matrix cells and is more robust under class imbalance (Guillaro et al., 2023)), and Average Precision (area under the Precision-Recall curve (Kwon and others, 2025)).
Instance-level detection. When ground truth is expressed as bounding boxes rather than pixel masks—as in Tampered-IC13, OSTF (Qu et al., 2025), and ReceiptForgery (Mesquita et al., 2023)—evaluation follows the object detection tradition. A predicted box is a true positive if its box-IoU with a ground-truth annotation meets a threshold (commonly 0.5 or the COCO standard of 0.5:0.05:0.95). Average Precision (AP) is then the area under the precision-recall curve over all confidence thresholds; mAP averages AP across forgery classes or across the IoU sweep. OSTF introduces a structured matrix of forgery-engine test-set configurations and reports the mean F1 (mF) across all 81 settings to capture cross-engine generalization. Instance-level metrics are fundamentally incompatible with pixel-level metrics unless boxes are filled to binary masks or masks are reduced to their bounding rectangle, each conversion introducing information loss.
Image-level detection. For the binary question of whether an image has been tampered at all, the standard metric is image-level AUC-ROC, used by TruFor (Guillaro et al., 2023), MVSS-Net (Chen et al., 2021), and all AI-generated image detectors (Wang et al., 2020; Ojha et al., 2023; Wang et al., 2023). CNNDetection and UnivFD instead report image-level Average Precision (mAP) because it is insensitive to the fraction of fake images in the test set. Identity document benchmarks (FantasyID (Korshunov et al., 2025), SIDTD (CVC Barcelona, 2023)) follow the ISO/IEC 30107-3 biometric standard, reporting APCER (attack error rate) and BPCER (bona fide error rate) rather than forensics metrics.
Implications for this benchmark. The fragmentation across metric conventions makes cross-paper comparison unreliable. DocForge-Bench standardizes on two complementary metrics applied uniformly to all 14 methods: Pixel-F1 at (deployment-relevant localization) and Pixel-AUC (calibration-independent ranking). Reporting both jointly is the key diagnostic: a high AUC with near-zero F1 is the calibration gap that is our central empirical finding, invisible under single-threshold protocols.
@@ 2.5 Existing Benchmarks and Surveys
Terminology note. Throughout this paper, zero-shot evaluation refers to applying any method with its published pretrained weights and no domain adaptation; we use out-of-distribution (OOD) to denote any (method, dataset) pair where the dataset was unseen during training.
@@ 3.1 Text-Region Tampering
We evaluate this threat across six complementary datasets spanning different document domains and imaging conditions: DocTamper (Qu et al., 2023a) provides large-scale synthetic diversity (170K images) across document layouts. T-SROIE (Zhang and others, 2022) captures receipt-context tampering with realistic OCR backgrounds. RealTextManipulation (Liao and others, 2022) provides authentic forgeries collected in the wild, without synthesis artifacts. Tampered-IC13 (Wang et al., 2022b) covers in-the-wild scene-text images (storefronts, signage) rather than document scans. FSTS-1.5k (Yu et al., 2025) provides a real-world evaluation derived from 16,750 human-annotated forgery instances, capturing the authentic distribution of text forgery parameters.
@@ 3.4 Scope and Coverage
Table 1 maps each dataset to the threat scenarios above and summarises evaluation coverage. Together, these seven datasets span four dimensions of variation critical for robust evaluation: (1) Realism: synthetic (DocTamper) vs. real-world (RealTextManipulation, ReceiptForgery, FSTS-1.5k); (2) Annotation: pixel-level masks vs. bounding boxes (ReceiptForgery, Tampered-IC13); (3) Document type: formal documents, receipts, scene text, and identity cards; (4) Evaluation scale: from 35 images (ReceiptForgery forged test set) to 1,488 (FSTS-1.5k full set); see the Eval. Images column in Table 1 for complete evaluation sizes.
We do not evaluate AI-generated document content (GANs, diffusion models), LLM-generated text detection, or physical print-scan-reprint attacks in the current release of DocForge-Bench. These represent important open challenges; AIGC-based document forgery is discussed as a critical open direction in Section 10, and the evaluation toolkit is designed to support such extensions without architectural changes.
@@ 4.1 Document-Specific Tampering Datasets
(Liao and others, 2022) provides 9,000 real-world text manipulation images collected from the internet, with pixel-level semantic segmentation masks. Unlike synthetically generated benchmarks, this dataset captures authentic forgeries with diverse backgrounds, text styles, and manipulation strategies. The official test split (3,197 images) contains 1,203 tampered images and 1,994 authentic images (“good_*” prefix, empty masks); we evaluate on the tampered-only subset (1,000 sampled after pre-filtering authentic images, which would otherwise yield NaN under our convention and be excluded from the dataset mean).
@@
(Mesquita et al., 2023) (L3i / ICDAR 2023 Competition) contains 988 receipt photographs (577 train / 193 val / 218 test), of which 163 across all splits are forged; the test split contains 35 forged images out of 218 (16% positive rate). Forgeries were produced with common image editors (GIMP, Paint) by replacing printed price or quantity fields. Annotations are VGG-format rectangular bounding boxes embedded in a CSV manifest; pixel masks are rasterized from bounding boxes for localization evaluation. Of the 218 test images, 183 are authentic (empty masks) and are excluded from pixel-level metric computation via the NaN convention; metrics are aggregated over the 35 forged images only. Due to the small positive test set (), per-method results on this dataset carry higher variance and should be interpreted with caution.
@@
(Anonymous, 2024) is a multi-label document tampering dataset containing approximately 30,200 images with five tampering categories: copy-move, splicing, text-generating, smearing, and erasing. Ground truth is provided as RGB color-coded masks where each channel indicates a distinct tampering type; binary evaluation maps any non-zero pixel to tampered. This any-channel binarization may inflate recall for methods sensitive to only one tampering type; method ranking on MixTamper should be interpreted with this simplification in mind. Images are uniformly cropped to pixels and sourced from the StaVer and SCUT-EnsExam corpora. We use the 6,817-image test split and sample 1,000 for tractable evaluation.
@@
| Dataset | Domain | Year | Dataset Size | Eval. Images | Forgery Types | GT Masks | Used In |
| DocTamper | Document | 2023 | 170K | 1,000 (sampled)† | Text replace/insert/delete | ✓ | Both |
| T-SROIE | Receipt | 2022 | 360 | 360 | Text tampering | ✓ | Both |
| RealTextManip. | Real-world | 2022 | 9K | 1,000 (sampled)† | Text manipulation | ✓ | Both |
| Tampered-IC13 | Scene text | 2015 | 233 | 188 tampered | Text replacement | BBox∗ | Both |
| Receipt Forgery | Receipt | 2023 | 988 | 35 forged | Price/qty forgery | BBox∗ | Both |
| MixTamper | Document | 2024 | 30K | 1,000 (sampled)† | Copy-move, splice, gen., smear, erase | ✓(RGB) | Both |
| FSTS-1.5k | Real-world | 2025 | 1,488 | 1,488 | Text manipulation | ✓ | Both |
| FantasyID | Identity doc. | 2025 | 6,571 | 2,773 tampered | Face swap, text replace | ✓ | Both |
| †Datasets with (sampled) draw 1,000 images at random (seed 42) from the full test split for tractable evaluation. | |||||||
@@ 5.2 Document-Specific Tampering Detection
All methods are evaluated with their official pretrained weights without fine-tuning. The general forensics methods (TruFor, ManTraNet, MVSS-Net, CAT-Net, PSCC-Net, IML-ViT, SAFIRE) were trained exclusively on natural photographic imagery—CASIA (Dong et al., 2013), FantasticReality (Kniaz et al., 2019), IMD2020 (Novozamsky et al., 2020), tampCOCO, RAISE, MS COCO, camera-model datasets (VISION, Dresden, KCMI), and proprietary synthetic databases. None saw any document images; all eight benchmark datasets are fully out-of-distribution for them. Among document-specific methods, DocTamper (model), DTD, FFDN, CAFTB-Net, and ADCD-Net were each trained on the DocTamperV1 training split (120,000 images)—their DocTamper test results are therefore in-domain, while all other datasets remain cross-domain. ASCFormer was trained on the RTM training split only (5,803 images); its RTM test results are in-domain. TIFDM was trained on a private, undisclosed corpus by the original authors (Dong et al., 2024); its DocTamper performance (F1=0.742) substantially exceeds a from-scratch DocTamper baseline, which may indicate in-domain training overlap, though this cannot be confirmed. All seven non-DocTamper datasets are definitively cross-domain for TIFDM.
@@ 6.1 Evaluation Protocol
Image-level detection (future extension).
Three datasets in DocForge-Bench—Tampered-IC13, ReceiptForgery, and FantasyID—include authentic (untampered) images alongside forged ones, enabling image-level binary detection evaluation. For each method, an image-level score can be derived by aggregating the predicted pixel map (e.g., maximum predicted score, or mean of the top-1% of pixels). Image-level AUROC on these three datasets would directly address the practical question of whether a document should be flagged for inspection. We defer this evaluation to a future extended version; the prediction infrastructure is in place and the analysis script is released alongside the benchmark toolkit.
@@ 6.2 Experiments
We compare seven document-specific methods (DocTamper, DTD, FFDN, CAFTB-Net, TIFDM, ASCFormer, ADCD-Net) against all seven general methods (TruFor, ManTraNet, MVSS-Net, CAT-Net, PSCC-Net, IML-ViT, SAFIRE) on all eight document datasets. For datasets with bounding-box annotations only (ReceiptForgery, Tampered-IC13), predicted masks are evaluated against rasterized ground-truth boxes. Authentic images (empty GT masks) are excluded from pixel-level metric aggregation via the NaN convention. This experiment directly measures the zero-shot advantage of domain-specific training and whether recent document forensics advances outperform both DocTamper and general forensics methods.
@@ 6.3 Implementation Details
All experiments are conducted using an NVIDIA GPU with at least 24 GB memory. Images are processed at each method’s native input resolution. For pixel-level evaluation, prediction masks are resized to match ground truth dimensions using bilinear interpolation. To maintain tractable evaluation on high-resolution datasets (T-SROIE receipt images can reach 49617016 px; FantasyID images average 32002000 px), both predictions and ground-truth masks are jointly downsampled to at most 2 megapixels using area interpolation before metric computation. This cap does not apply to images already within 2 MP. Evaluation scripts use scikit-learn for metric computation. For datasets exceeding the 1,000-image sample cap (DocTamper: 30K images; RealTextManipulation: 1,203 tampered images), images are selected using random.sample with seed 42, then sorted by filename for deterministic ordering. For ReceiptForgery, Pixel-IoU and Pixel-AUC are computed only over the 35 forged test images (16% of the 218-image test split); the remaining 183 authentic images produce NaN values under our convention and are excluded from the dataset mean.
@@
Methods vary in how they handle high-resolution inputs: TruFor processes images up to 40964096 px natively; MVSS-Net and PSCC-Net resize to their training resolution (512512 and 512512 respectively) before inference; IML-ViT pads/crops images to 10241024 for inference (only the top-left 10241024 region is evaluated on images larger than this; for FantasyID, an analysis of annotation bounding boxes shows that approximately 79% of forgery pixels fall outside this crop due to the distributed layout of identity card fields, so IML-ViT’s FantasyID score underestimates its true capability on full-card evaluation). IML-ViT’s FantasyID result is marked with § in Table 4. A tiled-inference variant (overlapping 10241024 crops with mask stitching) is a natural extension that would recover meaningful localization on high-resolution documents; we report the single-crop result for protocol consistency. This resolution heterogeneity constitutes a benchmark dimension in its own right: methods with larger effective receptive fields (TruFor at 40964096) are systematically advantaged on high-resolution datasets (FantasyID, T-SROIE) compared to methods resizing to 512512 (MVSS-Net, PSCC-Net). The complete evaluation pipeline, including configs, scripts, and metric computation, is released as open-source at https://github.com/BensonRen/document_forgery_benchmark.